Find the vector equation of the plane determined by the points (1,-2,1),(1,0,1) and (1,-1,1). Also check whether (1/2,1/2,1/2) lies on it.


1
Expert's answer
2021-02-16T15:46:27-0500

Solution.

If the coordinates of the three points "A(x_1;y_1;z_1)," "B(x_2;y_2;z_2)" and "C(x_3;y_3;z_3)" , which lie on the plane, the plane equation can be found by the following formula


"\begin{vmatrix}
x-x_1 & y-y_1 & z-z_1 \\
x_2-x_1 & y_2-y_1 & z_2-z_1 \\
 x_3-x_1 & y_3-y_1 & z_3-z_1 \\
 

\end{vmatrix}=0."

We will get


"\begin{vmatrix}
x-1 & y+2 & z-1 \\
0 & 2 & 0 \\
 0 & 1 & 0 \\
 

\end{vmatrix}=0,"

from here


"(x-1)•0+(y+2)•0+(z-1)•0=0,
\newline\text{or}\newline
0\cdot x+0\cdot y+0\cdot z+0=0."

"\vec{n}=\begin{pmatrix}
 0 \\
 0 \\
0
\end{pmatrix}" - normal vector.

The vector equation of the plane:


"<\vec{n},\vec{r}>=a,"

where "\vec{n}=\{A;B;C\}=\{0;0;0\}," "\vec{r}=\{x,y,z\}," and "a=-D=0."

So,


"\begin{pmatrix}
 0 & 0 & 0
\end{pmatrix}
×\begin{pmatrix}
 x\\
 y\\
z
\end{pmatrix}=0"

is the vector equation of the plane. It is not defined. The point (1/2;1/2;1/2) lies on it.